• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name X. Zheng

    Log entry time 02:36:58 on June17,2001

    Entry number 63471

    keyword=Target - EPR 37.7%

    2:17am, EPR during run 1609/20609, with beam on, did not see any wierd noise from photodiode. With #6 laser , oven 170, pol 3He cell position:

    Polarization up, down =

    0th order fit:
    37.73%, 37.03%

    1st order fit:
    37.68%, 37.03%

    Some adjustments:

    (1)Yesterday during 11:15am EPR, found lineshape was not symmetric (amplitude -10uV to +20uV. Had adjusted lock-in phase from 88.9 degree to 104.6 degree. Scanned by hand, got much better lineshape (more symmetric, -22 to +22uV).

    (2) EPR was noisy yesterday and the measurement has big error bar. So I adjusted PI gains before AFP flip during this measuremnt. 0-order fit shows:

    start point, end point, mean freq(Hz), sigma(Hz):
    5, 99, 11533867.11579, 1557.86575,
    120, 151, 11534121.00000, 853.31188,

    I did adjustment between point #99 and #151.